OpenStudy (anonymous):

Sabrina has a bag of 5 marbles that are identical in size. The bag contains 4 blue marbles and 1 red marble. If Sabrina randomly draws from the bag, replaces it then draws another marble, what is the probability that she will first draw a blue marble and then a red marble?

OpenStudy (anonymous):

see the process involves 2 steps, drawing the first ball and THEN the 2nd. first ball has to be a blue ball , so P(blue)=4/5. second ball has to be red so, P(red)=1/4. see the reduced denominator. this is because, since u've removed one blue ball already only 4 balls remain, hence 4. so answer is 4/5*1/4=1/5

OpenStudy (anonymous):

wy multiplication is, because u want both the events to happen . see carefully , in the question it says blue AND a red ball. so if and comes it is *, if Or comes it is + :)

OpenStudy (anonymous):

The answer key says 4/25... Any suggestions on why??

OpenStudy (anonymous):

oh! sorry it is , 4/25 only...it says he replaces the ball right, so wen u draw the second time it stilll remains 1/5 only...so 4/5*1/5=4/25. get it?

OpenStudy (anonymous):

yes=)